A 9900 N small car is moved by a 650N force. If the car starts from rest, then how fast is it traveling after 111 m? 


1
Expert's answer
2021-02-09T10:41:43-0500

a) Let's first find the mass of the car:


"m=\dfrac{W}{g}=\dfrac{9900\ N}{9.8\ \dfrac{m}{s^2}}=1010\ kg."

Then, we can find the acceleration of the car:


"a=\dfrac{F}{m}=\dfrac{650\ N}{1010\ kg}=0.64\ \dfrac{m}{s^2}."

Finally, we can find the velocity of the car after it traveled 111 meters:


"v_f^2=v_i^2+2ad,""v_f=\sqrt{v_i^2+2ad},""v_f=\sqrt{0+2\cdot0.64\ \dfrac{m}{s^2}\cdot111\ m}=12\ \dfrac{m}{s}."